In loading a delivery van, the driver pushed a 15 kg crate for
5 metres up a ramp. He had to push with a force of 132 N and
the vertical height gained was 1.3 m. What was the efficiency of
pushing this crate up the ramp onto the back of the van?
1
Expert's answer
2020-12-03T06:58:48-0500

Answer

Length of ramp

"=\sqrt{5^2+1.3^2}=5.17m"

Net work done is equal to change in potential energy

"W=U=mgh"

"=15\times10\times1.3"

"W=195J"

Total work done

"W=Fd=132\times1.57"

"W=682.44J"

Efficiency is given by

"\eta=\frac{195}{682.44}\times100=29.12\%"